A small stone of mass, is thrown vertically upward with initial speed ,u. If the air resistance at speed is mkv , where v is a positive constant. Show that the stone returns to its starting point with speed given by v=u/(( 1+ku .


1
Expert's answer
2020-01-20T05:18:48-0500

As per the question,

Resistivity force "F_r=-mkv"

And force due to gravity "F_g=-mg"

So, net force "ma= F_g+F_r"

"\Rightarrow ma=-(mg+mkv)"

we know that "a=\dfrac{dv}{dt}=\dfrac{dv}{dt}\times \dfrac{dx}{dx}=v\dfrac{dv}{dx}"

"\Rightarrow ma=-m(g+kv)"

"\Rightarrow v\dfrac{dv}{dx}=-(g+kv)"

"\Rightarrow \int_u^v \dfrac{vdv}{g+kv}=\int_0^x dx"

At the final condition, net displacement =0

(There is mistypes data in the question, either "mkv^2" or the answer given will be something else)